You have a pizza; if one person eats 4/9 of the pizza and another eats 1/12 of the pizza, how much of the pizza remains. Write the answer as a fraction in lowest terms, with appropriate units.
Answer by Boreal(15235)

Amount eaten is (4/9)+(1/12)
Common denominator is 36
(16/36)+(3/36)=19/36 eaten
That leaves 1-(19/36)=17/36 of a pizza